# Revisiting the Thickness of the Air-Water Interface from Two Extremes of Interface Hydrogen Bond Dynamics

### v2

@@ III Results and Discussions
In this section, we implement our two-extreme approach to two distinct system properties: HB population auto-correlation functions and HB reaction rate constants. We then determine the thickness of the air-water interface based on each of these properties.
III.1 HB population auto-correlation functions
@@
Figure 2: The $c(t)$ and $c^{(\text{s})}(t)$ for interface H-bonds with different thickness $d$ obtained from the ADH criterion of HB definition for (a) scenario 1 (LC); (b) scenario 2 (IHB). Two features can be found: (i). As $d$ increases, both $c(t)$ and $c^{\rm(s)}(t)$ eventually approach a stable function. (ii). The decay rate of $c^{(\text{s})}(t)$ in scenario 2 surpasses $c(t)$ in scenario 1. This is visually represented by two dashed directed line segments, positioned identically to the graphs.
@@
Figure 2 illustrates the dynamic evolution of $c(t)$ and $c^{\rm(s)}(t)$ for various values of $d$. When $d$ is equal to or greater than 4.0 (Å), the time evolution curves of $c(t)$ and $c^{\rm(s)}(t)$ become largely invariant to further increases in $d$. This shows that one can calculate the thickness of the interface under different conditions with the help of the dependence of the correlation function on $d$. Analyzing both scenarios, it is evident that the decay rate of the correlation function $c^{(\text{s})}(t)$ in scenario 2 (IHB) surpasses that in scenario 1 (LC).
@@
Figure  3 shows the $d$-dependence of the correlation functions at three reference time points $t^{*}=1,2,5$ (ps). It shows the $d$-dependence clearly from another angle: we selected the three reference time intervals on the $t$-axis in Fig. 2 and recorded the values of the correlation functions, $c(t)$ and $c^{\rm(s)}(t)$, at these time intervals for each $d$. Comparing $c(t)$ and $c^{(\text{s})}(t)$ in Fig. 3, $c(t)$ is always slightly larger than $c^{(\text{s})}(t)$ for the same $d$.
In Fig. 3, both $c(t^{*})$ and $c^{\rm(s)}(t^{*})$ increase as $d$ increases, and eventually their change rates approach 0 when $d$ is large enough. Since $c(t)$ and $c^{(\text{s})}(t)$ are the upper and lower bounds of the real interface correlation functions $c^{r}(t^{*})$ respectively. It logically follows that the $d$-dependence of $c^{r}(t^{*})$ has the same changing trend. Then we have the equations ${\rm d}C/{\rm d}d=0$, for $c(t^{*})$ and $c^{\rm(s)}(t^{*})$, with respect to $d$. For $C=c(t^{*})$ and $C=c^{\rm(s)}(t^{*})$, we can find solutions $d_{f1}$ and $d_{f2}$, respectively. Here $d_{f1}$ and $d_{f2}$ are the thickness of the air-water interface obtained from the LC method and the IHB method, respectively. Then $d_{f}=(d_{f1}+d_{f2})/2$ is obtained as the thickness of the interface for a given $t^{*}$.
@@
Figure 3: The $d$-dependence of the correlation functions at three reference time points $t^{*}=1,2,5$ (ps), revealing key insights into the dynamics of the air-water interface: (i). As $d$ increases, both $c(t^{*})$ and $c^{\rm(s)}(t^{*})$ show an upward trend, with their rates of change gradually approaching 0. (ii). For each $t^{*}$, $c(t)$ is always slightly larger than $c^{(s)}(t)$ for the same $d$. (iii). The thickness $d_{f}=4.0$ (Å) of the air-water interface is obtained from the $d$-dependence of $c(t^{*})$ and $c^{\rm(s)}(t^{*})$, obtained by the LC method (dot-dashed lines) and the IHB method (solid lines), respectively.
@@
The correlations $c(t^{*})$ and $c^{\rm(s)}(t^{*})$ respectively describe the relaxation characteristics of HB dynamics within the interface. For both the LC and IHB methods, the values of $c(t^{*})$ and $c^{\rm(s)}(t^{*})$ no longer change significantly with the increase of $d$, when $d$ increases to 4.0 (Å), ie., $d_{f1}\approx d_{f2}=4.0$ (Å). Therefore, we obtain the thickness of the air-water interface in the simulations, i.e., $d_{f}=4.0$ (Å). Note that the ADH criterion of H-bonds is used here. Such an observation is in harmony with the findings based on the AHD criterion, as elaborated in Appendix A.
@@ III.2 HB reaction rate constants
Figure 4: The breaking HB reaction rate constants $k_{\rm LC}$ and $k_{\rm IHB}$ obtained by the IHB and LC methods respectively. (i). Both constants, $k_{\rm LC}$ and $k_{\rm IHB}$, decrease monotonically to the HB breaking rate $k_{\rm bulk}$ for the bulk water as $d$ increases. (ii). $k_{\rm LC}$ is always smaller than $k_{\rm IHB}$ for the same $d$. (iii). The thickness $d^{\prime}_{f}=4.0$ (Å) of the air-water interface is obtained from the $d$-dependence of the rate constant $k_{\rm IHB}$ and $k_{\rm LC}$, obtained by the IHB method and the LC method, respectively.
We further examined how the reaction rate constants of H-bonds at the air-water interface vary with $d$. For further details regarding the HB reaction rate constants, please refer to Appendix B.
In Fig. 4, we compare the breaking HB reaction rate constantsLuzar and Chandler (1996a); Luzar (2000), $k_{\rm LC}$ and $k_{\rm IHB}$, obtained by the IHB and LC methods, respectively. We found that both constants, $k_{\rm LC}$ and $k_{\rm IHB}$, decrease monotonically as $d$ increases. When $d$ is large, both rate constants $k_{\rm IHB}$ and $k_{\rm LC}$ also no longer change with $d$. In addition, the HB breaking rate $k_{\rm bulk}$ for the bulk water is also drawn as a reference. As $d$ increases, both $k_{\rm LC}$ and $k_{\rm IHB}$ gradually close to the rates $k_{\rm bulk}$. It also shows that the HB beaking rate constant $k_{\rm IHB}$ is relatively larger than $k_{\rm LC}$. This difference is related to the definitions of HB populations, $h(t)$ and $h^{\text{(s)}}(t)$. The definition of $h^{(\text{s})}(t)$ makes the HB break rate at the interface increase. The LC method, which is based on Luzar-Chandler HB population $h(t)$, retains the original rate constant of H-bonds, but it may include the contribution of bulk water molecules to the rate constant.
Similar to the HB auto-correlation functions, we have the equations ${\rm d}K/{\rm d}d=0$, for $k_{\rm IHB}$ and $k_{\rm LC}$, with respect to $d$. For $K=k_{\rm IHB}$ and $K=k_{\rm LC}$, we can find solutions $d^{\prime}_{f1}$ and $d^{\prime}_{f2}$, respectively. Here $d^{\prime}_{f1}$ and $d^{\prime}_{f2}$ are also the HB reaction rate-based thickness of the interface obtained from the LC and IHB method, respectively. Then we obtain the thickness of the real air-water interface as $d^{\prime}_{f}=(d^{\prime}_{f1},d^{\prime}_{f2})/2\approx 4.0$ (Å). This result is supported by our calculation of the $d$-dependence of the HB reforming rate constants, ie., $k^{\prime}_{\rm LC}$ and $k^{\prime}_{\rm IHB}$. For more details, please refer to Fig. B7 in Appendix B).
@@ III.3 Orientational relaxation of the OH stretch and other supports
To verify our conclusion on the thickness of the air-water interface from the interfacial HB dynamics, we calculated the orientational relaxation of the OH stretch from the perspective of anisotropy decay. As shown in Fig. 5, the values of the orientational correlation function $C_{2}(t)$ at four reference times $t^{*}=1,2,5$ (ps) are plotted. It can be seen that when the interface thickness $d$ is small, the OH orientational relaxation time $\tau_{2}$ of the $C_{2}(t)$ increases linearly with $d$. The $\tau_{2}$ for the neat water surface is $1.4$ to $2.0$ (ps), with std. error $\delta d=0.1$ (ps), which is consistent with the vibrational relaxation time of the free OH for the neat water ichi Inoue et al. (2020).
@@
Similar to the HB population correlation functions $c(t)$ and $c^{\rm(s)}(t)$ and the HB reaction rates, $k_{\rm LC}$ and $k^{\prime}_{\rm LC}$, and $k_{\rm IHB}$ and $k^{\prime}_{\rm IHB}$, as $d$ increases, $C_{2}(t^{*})$ and $\tau_{2}$ converges to a fixed value, which characterizes the decay time of the orientation relaxation process of OH bonds in bulk water. From the convergence trend of $\tau_{2}$ in Fig. 5(b), we find that at the interface with a thickness greater than 4.0 Å, the OH orientation relaxation of the air-water interface is no longer different from bulk water. Therefore, we arrive at a consistent conclusion on the issue of estimating the thickness of the air-water interface, from the perspective of HB dynamics and OH reorientation relaxation. Furthermore, the interface thickness aligns with values obtained through experimental measurements and AIMD simulations, as detailed in Table 1.
Figure 5: The $d$-dependence of (a) $C_{2}(t^{*})$ and (b) $\tau_{2}(t^{*})$ for water molecules at the air-water interface with $t^{*}=1,2,5$ (ps), respectively. As $d$ increases to $d=4$ (Å), $C_{2}(t^{*})$ and $\tau_{2}$ converge to a fixed value, respectively.
@@
Table 1: The air-water interface thickness obtained by experiments and computer simulations ("-" denotes without using "10–90" thickness).
Method Thickness Notation $T$ (K) $d$ (Å)
Ellipsometry (Rayleigh) - 293.15 3.0
Ellipsometry (Raman and Ramdas (1927)) - 293 5.0
Ellipsometry (McBain, Bacon, and Bruce (1939)) - 293 $\geq$2.26
Ellipsometry (Kinosita and Yokota (1965)) - 293 7.1
X-Ray Reflectivity(Braslau et al. (1988)) - 298 3.24$\pm$ 0.05
MD (TIP4P/FQ, Liu et alLiu, Harder, and Berne (2005)) 10–90 298.15 3.5
MD (TIP4P-POL2, Kuo et alKuo et al. (2006)) 10–90 298 3.9
MD (SPC-FW, Wick et alWick et al. (2007)) 10–90 298 3.2
MD (D-C, Wick et alWick et al. (2007)) 10–90 298 3.2
MD (TIP4P, Wick et alWick et al. (2007)) 10–90 298 3.4
MD (SPC/E, Fan et alFan et al. (2009)) 10–90 300 6.0
MD ( SPC/E, Willard and ChandlerWillard and Chandler (2010)) 10–90 298 8.0
MD (SPC/E, Vila Verde et al. Vila Verde, Bolhuis, and Campen (2012)) 10–90 300 3.27
MD (TIP4P/2005, Vila Verde et al. Vila Verde, Bolhuis, and Campen (2012)) 10–90 300 3.05
CPMD/BLYP (Kuo et alKuo et al. (2006)) 10–90 300 1.7
BOMD/BLYP-D (Kühne et al. Kühne et al. (2011)) 10–90 300 3.8
BOMD/BLYP-D2 (Pezzotti et alPezzotti, Galimberti, and Gaigeot (2017) ) 10–90 315 3.5
BOMD/BLYP-D3 (Jeon el al. Jeon et al. (2017)) 10–90 298 $\sim$3.0
BOMD/BLYP-D3/LC&IHB (this work) - 300 4.0
@@ IV Conclusions
In this study, we have developed a "two-extremes" approach to investigate the HB dynamics at the air-water interface and to determine the interface’s thickness. One extreme scenario underestimates the HB breaking rate constant, while the other overestimates it, implying that each scenario provides only a partial insight into the interfacial HB dynamics. Subsequently, based on the DFTMD simulations, we have applied our approach to two distinct system properties: HB relaxation and HB reaction rate constants at the air-water interface. Our results across all these properties indicate that the predictions from both extreme scenarios converge as the thickness of the air-water interface increases to 4.0 Å. Thus, we have reason to believe the thickness, which falls between these two extremes, converges at this critical value.
This work introduces an approach to complement existing ones for investigating the air-water interface from a fresh perspective. Through HB dynamics of the air-water interface, interfacial properties, such as thickness in this case, can be obtained through a method analogous to the squeeze theorem. Beyond the scope of HB dynamics, this approach can be extended to other properties like molecular orientation distribution Rao, Tao, and Wang (2003); Gan et al. (2006b) and SFG spectrum Nihonyanagi et al. (2011, 2015); Tang et al. (2020), as well as other solution interfacial surfaces where statistical properties of the interface and bulk phase differ significantly. Looking ahead, our study could also inspire researchers to examine the hydration shells of ions, and then study the effects of ions through HB dynamics.

### v3

@@ III Results and Discussions
In this section, we apply our two-extreme approach to two system properties: HB population autocorrelation functions and HB reaction rate constants. We then determine the thickness of the air-water interface based on each of these properties. The computational details for DFTMD and DeePMD simulations are available in the Supporting Information. The results discussed in the following sections are derived from the trajectory data of the system containing 512 water molecules, modeled by MB-pol potential. The finite size effects of the main properties discussed in this article are in Supporting Information. Similar analyses for DFTMD simulations with 128 water molecules, and other system sizes in DeePMD simulations using MB-pol, are also provided in the Supporting Information.
III.1 HB Population Autocorrelation Functions
@@
Figure 2: Autocorrelation functions $c(t)$ and $c^{(\text{s})}(t)$ for interface H-bonds with different thickness $d$ for (a) Scenario 1 (LC) and (b) Scenario 2 (IHB). Two notable features emerge: (i) As $d$ increases, both $c(t)$ and $c^{\rm(s)}(t)$ eventually approach a stable function. (ii) Decay rate of $c^{(\text{s})}(t)$ in Scenario 2 is greater than that of $c(t)$ in Scenario 1. This behavior is visually represented by two dashed directed line segments, positioned identically to the graphs. Functions $c(t)$ and $c^{(\text{s})}(t)$ on a log-log scale are also plotted in Supporting Information.
@@
Figure 2 illustrates the dynamic evolution of $c(t)$ and $c^{\rm(s)}(t)$ for various values of distance $d$. When $d$ is 4 (Å) or greater, $c(t)$ and $c^{\rm(s)}(t)$ become largely invariant to further increases in $d$. This indicates that one can determine the thickness of the interface under different conditions by examining how the correlation function depends on distance $d$. Analysis of both scenarios reveals that the decay rate of the correlation function $c^{(\text{s})}(t)$ in Scenario 2 (IHB) surpasses that in Scenario 1 (LC).
@@
Figure  3 displays the $d$-dependence of the correlation functions at three reference time points $t^{*}=1,2,5$ (ps). This figure provides a different perspective on $d$-dependence: by selecting three reference time intervals on the $t$-axis in Figure 2, the values of the correlation functions, $c(t)$ and $c^{\rm(s)}(t)$, at these time intervals for each $d$ were recorded. Comparing $c(t)$ and $c^{(\text{s})}(t)$ in Figure 3, $c(t)$ in Scenario 1 is always slightly larger than $c^{(\text{s})}(t)$ in Scenario 2 for the same $d$.
@@
Figure 3: Dependence of the correlation functions on the distance $d$ at three reference time points $t^{*}=1,2,5$ (ps) provides key insights into the dynamics of the air-water interface: (i) As $d$ increases, both $c(t^{*})$ and $c^{\rm(s)}(t^{*})$ exhibit an upward trend, with their rates of change gradually approaching 0. (ii) At each $t^{*}$, $c(t)$ is consistently slightly larger than $c^{\rm(s)}(t)$ for the same $d$. (iii) Air-water interface thickness $d_{f}=4$ (Å) is determined from the $d$-dependence of $c(t^{*})$ and $c^{\rm(s)}(t^{*})$, as calculated using the LC method (dot-dashed lines) and the IHB method (solid lines), respectively.
In Figure 3, both $c(t^{*})$ and $c^{\rm(s)}(t^{*})$ increase as $d$ increases, eventually reaching a point where their change rates approach 0 for sufficiently large values of $d$. Since $c(t)$ and $c^{(\text{s})}(t)$ serve as the upper and lower bounds of the real interface correlation function $c^{r}(t^{*})$ respectively, it logically follows that the $d$-dependence of $c^{r}(t^{*})$ exhibits the same trend. Consequently, we arrive at the equations ${\rm d}C/{\rm d}d=0$ for $c(t^{*})$ and $c^{\rm(s)}(t^{*})$ with respect to $d$. Solving for $C=c(t^{*})$ and $C=c^{\rm(s)}(t^{*})$ yields solutions $d_{f1}$ and $d_{f2}$, respectively. Here $d_{f1}$ and $d_{f2}$ represent the thickness of the air-water interface as determined by the LC method and the IHB method, respectively. The average value $d_{f}=(d_{f1}+d_{f2})/2$ is then obtained as the thickness of the interface for a given $t^{*}$.
@@
The correlations $c(t^{*})$ and $c^{\rm(s)}(t^{*})$ respectively describe the relaxation characteristics of HB dynamics within the air-water interface. For both the LC and IHB methods, the values of $c(t^{*})$ and $c^{\rm(s)}(t^{*})$ cease to change significantly when $d$ reaches 4 (Å), i.e., $d_{f1}\approx d_{f2}=4$ (Å). Thus, we determine the thickness of the air-water interface in the simulations to be $d_{f}=4$ (Å).
@@ III.2 HB Reaction Rate Constants
We further examined how the reaction rate constants of H-bonds at the air-water interface vary with $d$. For further details regarding the HB reaction rate constants, please refer to Section 2 in Supporting Information.
In Figure 4, we compare the breaking HB reaction rate constantsLuzar and Chandler (1996); Luzar (2000), $k_{\rm LC}$ and $k_{\rm IHB}$, obtained by the LC and IHB methods, respectively. We found that both constants, $k_{\rm LC}$ and $k_{\rm IHB}$, decrease monotonically as $d$ increases. When $d$ is large, both rate constants $k_{\rm IHB}$ and $k_{\rm LC}$ also no longer change with $d$. It also shows that the HB breaking rate constant $k_{\rm IHB}$ is relatively larger than $k_{\rm LC}$. This difference is related to the definitions of HB populations, $h(t)$ and $h^{\text{(s)}}(t)$. The definition of $h^{(\text{s})}(t)$ leads to an increased HB break rate at the interface. The LC method retains the original rate constant of H-bonds but may include contributions from bulk water molecules.
Figure 4: Breaking HB reaction rate constants $k_{\rm LC}$ and $k_{\rm IHB}$ for the air-water interface simulated with 512 water molecules, obtained by the LC and IHB methods respectively. (i) Both constants, $k_{\rm LC}$ and $k_{\rm IHB}$, decrease monotonically to the HB breaking rate $k_{\rm bulk}$ for the bulk water as $d$ increases. (ii) $k_{\rm LC}$ is always smaller than $k_{\rm IHB}$ for the same $d$. (iii) Thickness $d^{\prime}_{f}=4$ (Å) of the air-water interface is obtained from the $d$-dependence of the rate constant $k_{\rm LC}$ and $k_{\rm IHB}$, obtained by the LC method and the IHB method, respectively. (Consistent results for systems with 125 and 216 water molecules, please refer to the Supporting Information.)
Similar to the HB autocorrelation functions, we have the equations ${\rm d}K/{\rm d}d=0$, for $k_{\rm IHB}$ and $k_{\rm LC}$, with respect to $d$. For $K=k_{\rm IHB}$ and $K=k_{\rm LC}$, we identify solutions $d^{\prime}_{f1}$ and $d^{\prime}_{f2}$, respectively. Here $d^{\prime}_{f1}$ and $d^{\prime}_{f2}$ represent the HB reaction rate-based thickness of the interface obtained from the LC and IHB method, respectively. We then calculate the thickness of the real air-water interface as $d^{\prime}_{f}=(d^{\prime}_{f1}+d^{\prime}_{f2})/2\approx 4$ (Å). This result is supported by our calculation of the $d$-dependence of the HB reforming rate constants, namely, $k^{\prime}_{\rm LC}$ and $k^{\prime}_{\rm IHB}$. For more details, please refer to Figure S7 in Supporting Information.
@@ III.3 Orientational Relaxation of the OH Stretch and Other Supports
To verify our conclusion on the thickness of the air-water interface from the interfacial HB dynamics, we performed calculations on the orientational relaxation of the OH stretch from the perspective of anisotropy decayMoilanen et al. (2008).
Figure 5: Orientation correlation function $C_{2}(t)$ for the air-water interface. (a) The $C_{2}(t)$ for $d=1,4,6$ (Å) and the biexponential decay function $C_{2}(t)=A_{1}\exp(t/\tau_{1})+A_{2}\exp(t/\tau_{2})$. $d$-dependence of (b) $C_{2}(t^{*})$ for water molecules at the air-water interface with $t^{*}=1,2,5$ (ps); (c) relaxation time ($\tau_{1}$ and $\tau_{2}$), and (d) the amplitude ($A_{1}$ and $A_{2}$). Dotted line is a horizontal line for easier viewing. As $d$ increases to $d=4$ (Å), $C_{2}(t^{*})$ and $\tau_{i}$, $A_{i}$ all converge to a fixed value, respectively. (We also obtained consistent results for systems with 125 and 216 water molecules, please refer to Supporting Information.)
As shown in Figure 5a, the orientation correlation function fits well with a bi-exponential decay function. The two relaxation rates ($1/\tau_{1}$ and $1/\tau_{2}$) differ; the larger relaxation time $\tau_{2}$ increases with the increase of $d$, while the smaller one, $\tau_{1}$, remains relatively unchanged with variations in $d$. In Figure 5b, the values of the orientational correlation function $C_{2}(t)=\langle P_{2}(\hat{u}(0)\hat{u}(t))\rangle$ at three reference time $t^{*}=1,2,5$ (ps) are plotted. At all these times, $C(t^{*})$ increases with increasing $d$ and no longer increases significantly for $d\geq 4$ (Å).
@@
Similar to the earlier results on HB population correlation functions and the HB reaction rates, as $d$ increases, $C_{2}(t^{*})$ and associated relaxation times converge to a fixed value, respectively, which characterizes the decay time of the orientation relaxation process of OH bonds. From the convergence trend of the relaxation times and corresponding amplitudes in Figure 5c,d, we find that at the interface with a thickness greater than 4 Å, the OH orientation relaxation of the air-water interface is no longer different from bulk water.
Experimentally, the information on the interface molecules can be obtained from the SFG spectrum. The vibration relaxation time of the interface water molecules can be obtained based on the SFG spectrum using analysis techniques such as Singular Value Decomposition (SVD)ichi Inoue et al. (2020). We also defined the interface molecule orientation relaxation time from the $C_{2}(t)$ correlation of the interface water molecules. However, due to the arbitrariness of the specific form of defining the relaxation time, the results we obtained cannot be directly and accurately compared with the experimental results. Because of the current huge challenges in directly measuring water, we compared the data obtained from interfacial HB dynamics, interfacial OH orientation relaxation dynamics, and AIMD simulations. Our results from interfacial HB dynamics are consistent with the results from $C_{2}(t)$ and density profile from AIMD simulations of the air-water interface. Furthermore, the interface thickness aligns with values obtained through experimental measurements and AIMD simulations, as detailed in Table 1. Therefore, we arrive at a consistent conclusion on the issue of estimating the thickness of the air-water interface, from the perspective of HB dynamics and OH reorientation relaxation.
@@
Table 1: Air-Water Interface Thickness Obtained by Experiments and Computer Simulations.
Methods $T$ (K) $d$ (Å)
ellipsometry (Rayleigh) 293.15 3.0
ellipsometry (Raman and Ramdas (1927)) 293 5.0
ellipsometry (McBain, Bacon, and Bruce (1939)) 293 $\geq$2.26
ellipsometry (Kinosita and Yokota (1965)) 293 7.1
X-ray reflectivity(Braslau et al. (1988)) 298 3.24$\pm$ 0.05
DeePMD/MB-pol/Free OH (this work) 300 5.0
BOMD/BLYP-D3/LC&IHB (this work) 300 4.0
DeePMD/MB-pol/LC&IHB (this work) 300 4.0
There is one more aspect worth noting when studying the air-water interface based on the aforementioned definition of H-bonds. Some complications arise in interpreting vibrational SFG spectroscopy results Tang et al. (2017). An alternative perspective focusing on nonbonded OH groups better establishes a direct correlation between free OH and the 3700 cm-1 peak, typically attributed to free OH groups Tang et al. (2017). Considering the dynamics of free OH groups, the thickness of the air-water interface is also estimated from simulations. Details on the methods and results can be found in Supporting Information. From the $d$-dependence of the interfacial free OH correlation function, we determined an interface thickness of approximately 5 Å. As shown in Table 1, this value slightly deviates from the main results presented in this paper, highlighting that the measured thickness of the air-water interface varies depending on the properties of interest. Thus, to accurately determine the water interface thickness, it is essential to integrate findings from these varied perspectives.
@@ IV Conclusions
In this study, we have developed a two-extremes approach to investigate the HB dynamics at the air-water interface and to determine the interface’s thickness. One extreme scenario underestimates the HB breaking rate constant, while the other overestimates it, implying that each scenario provides only a partial insight into the interfacial HB dynamics. Subsequently, based on DFTMD and DeePMD simulations, we have applied our approach to two distinct system properties: HB relaxation and HB reaction rate constants at the air-water interface. Our results across both properties indicate that the predictions from both extreme scenarios converge as the thickness of the air-water interface increases to 4 Å. Thus, we have reason to believe the thickness, which falls between these two extremes, converges at this critical value.
This work introduces an approach to complement existing ones for investigating the air-water interface from a fresh perspective. Through HB dynamics of the air-water interface, interfacial properties, such as thickness in this case, can be obtained through a method analogous to the squeeze theorem. Beyond the scope of HB dynamics, this approach can be extended to other properties like molecular orientation distribution Rao, Tao, and Wang (2003); Gan et al. (2006b), free OH dynamics Tang et al. (2017), and SFG spectrum Nihonyanagi et al. (2011, 2015); Tang et al. (2020), or other systems like solution interfacial surfaces where statistical properties of the interface and bulk phase differ significantly. Looking ahead, this study could inspire further research into ions’ hydration shells by examining ions’ effects through HB dynamics.
It is important to note that our current approach focuses on probing the local environment by analyzing water pairs within HB dynamics, without considering the collective behaviors of water molecules in the HB networks. However, the collective dynamics of many water molecules have been increasingly discussed in recent research, demonstrating a close relationship with observable properties such as dielectric spectroscopy and time-dependent vibrational spectroscopy.Popov et al. (2016); Fecko et al. (2003, 2005); Hölzl, Forbert, and Marx (2021); Offei-Danso et al. (2023); Malosso et al. (2024) We hope our study will inspire further innovative ideas on related topics by incorporating considerations of the collective nature of water.
Notes
The authors declare no competing financial interest.
Code availability
The codes utilized in this study are publicly accessible on GitHub at https://github.com/hg08/hb_ihb.